# Linear instability of hairy black holes in Horndeski theory

### v1

@@ II Background fields and odd-parity gravitational perturbation
with $\kappa\equiv\sqrt{8\pi}$. For the 4-current to vanish at the infinity and the finiteness of energy, we require $\alpha_{21}=\alpha_{31}=0$ and $w_{2}=3w_{4}=3/2$ Bergliaffa2021 . The comma in the subscript stands for the derivative with respect to the quantity right after that, $R$ is the Ricci scalar, $\nabla$ denotes the covariant derivative operator, while the operator $\Box$ is defined as $\Box\equiv\nabla^{\mu}\nabla_{\mu}$.
@@ IV Conclusions
First of all, certain constraints to the coupling parameters are identified from the no-ghost condition [cf., (III)]. Based on that, the following calculations can be simplified. This enables us to calculate and analyze the radial Laplacian stability condition within the valid phase space of (reduced) coupling parameters (cf., Fig.1). In Fig.1 we learn that such a stability condition requires additional constraints to the coupling parameters. More importantly, this stability condition tends to be broken in the neighborhood of event horizon $r=1$. Indeed, by using the asymptotic expansion of the criterion ${\hat{c}}_{r}^{2}$ [cf., (3.2)], it is clear that this condition will not be preserved anymore as $r\to 1$ [although we also notice that such a condition can be satisfied right at the point $r=1$, since we precisely have ${\hat{c}}_{r}^{2}(r=1)=0$] no matter how the coupling parameters are chosen. We thus conclude that the set of background solutions (II) are not stable and needs to be excluded from the valid solutions, which is consistent with the predictions given in Minamitsuji2022 .
Notice that, setting those coupling parameters to zero (as in the GR limit) in the final criterion or discriminant [cf., (3.2)] will not always simply bring us back to that of GR, different from what we have seen at the Lagrangian level. In Tsujikawa21 ; Kobayashi2014 we also observed similar phenomena. This implies that (in)stability of a theory could be a comprehensive effect and it is not necessary for the criterion to behave like a perturbation to the stable GR case. It is also interesting to mention that, as we see from the Fig.1, the most stringent constraint to the coupling parameters are given by the analysis within the neighborhood of $r=1$ (instead of the other positions that far from the event horizon). That is similar to the situation shown in Kase2018 .

### v2

@@ II Background fields and odd-parity gravitational perturbation
with $\kappa\equiv\sqrt{8\pi}$. For the 4-current to vanish at infinity and the finiteness of energy, we require $\alpha_{21}=\alpha_{31}=0$ and $w_{2}=3w_{4}=3/2$ Bergliaffa2021 . The comma in the subscript stands for the derivative with respect to the quantity right after that, $R$ is the Ricci scalar, $\nabla$ denotes the covariant derivative operator, while the operator $\Box$ is defined as $\Box\equiv\nabla^{\mu}\nabla_{\mu}$.
@@ IV Conclusions
First of all, certain constraints to the coupling parameters are identified from the no-ghost condition [cf., (III)]. Based on that, the following calculations can be simplified. This enables us to calculate and analyze the radial Laplacian stability condition within the valid phase space of (reduced) coupling parameters (cf., Fig.1). In Fig.1 we learn that such a stability condition requires additional constraints to the coupling parameters. More importantly, this stability condition tends to be broken in the neighborhood of event horizon $r=1$. Indeed, by using the asymptotic expansion of the criterion ${\hat{c}}_{r}^{2}$ [cf., (3.2)], it is clear that this condition will not be preserved anymore as $r\to 1$ [although we also notice that such a condition can be satisfied right at the point $r=1$, since we precisely have ${\hat{c}}_{r}^{2}(r=1)=0$] no matter how the (physical) coupling parameters are chosen. We thus conclude that the set of background solutions (II) are not stable and needs to be excluded from the valid solutions, which is consistent with the predictions given in Minamitsuji2022 .
As observed from different occasions, setting those coupling parameters to zero (as in the GR limit) in the final criterion or discriminant [e.g., the (3.2) in this paper] will not always simply bring us back to that of GR, different from what we have seen at the Lagrangian level. In Tsujikawa21 ; Kobayashi2014 we also observed similar phenomena. This implies that (in)stability of a theory could be a comprehensive effect and it is not necessary for the criterion to behave like a perturbation to the stable GR case. It is also interesting to mention that, as we see from the Fig.1, the most stringent constraint to the coupling parameters are given by the analysis within the neighborhood of $r=1$ (instead of the other positions that far from the event horizon). That is similar to the situation shown in Kase2018 .